QUESTION 10 Komiko Tanaka invests $13,000 in LymaBean, Inc. LymaBean does not pay any dividends. Komiko projects that her investment will generate a 11 percent before-tax rate of return. She plans to invest for the long term. How much cash will Komiko retain, after-taxes, if she holds the investment for 6 years and then she sells it when the long-term capital gains rate is 15 percent?

QUESTION 11 Komiko Tanaka invests $13,000 in LymaBean, Inc. LymaBean does not pay any dividends. Komiko projects that her investment will generate a 11 percent before-tax rate of return. She plans to invest for the long term. If she holds the investment for 6 years and then she sells it when the long-term capital gains rate is 15 percent, what is her after-tax rate of return?
Answer #1
Question 10
Calculation of Capital gain and cash retained after tax at the end of 6th year
Particulars Calculation Amount
a Value of Investment after 6 years = 13000*(1+0.11)^6 = $ 24,315.39
b Less : Purchase cost of Investment $ 13,000.00
c Long term Capital gain a-b $ 11,315.39
d Tax on Long term Capital Gain at 15% d x 15% $   1,697.31
e Cash Retained after tax a-d $ 22,618.08
Question 11
Calculation of after tax rate of return
a Investment proceeds received after tax $ 22,618.08
b Initial Investment Amount $ 13,000.00
c So we can say that                                                                     13000*(1+x)^6 = 22618.08
13000*(1+x^6)= 22618.08
1+x^6=22618.08/13000
1+x^6= 1.73985231
x ^6 0.73985231
So x                0.12